Accelerator Update - August 4-August 6
The integrated luminosity for last week was 6.0 pb-1.
Operations put the TeV into shot setup at 9:04 AM.
A Booster expert began work on a power supply (ORBMP) at 12:41 PM. He found a small internal LCW leak. This halted stacking. He had the supply back in operation by 1:37 PM.
At 1:57 PM, the Main Injector (MI) suffered a vacuum pump trip and valve closure.
A MI expert began a slipstacking study at 3 PM. She finished her study at 4:34 PM.
Stacking stopped at 8:21 PM. Operators investigated and discovered a bad Pbar crate power supply. They changed the power supply and stacking resumed at 9:13 PM.
At 8:31 PM, at Linac RF station (LRF2) tripped off. It reset okay, but operators reported hearing a loud humming noise. A Linac expert suggested a fix when an appropriate moment occurred.
Tuesday August 5
TeV experts began an end-of-store study at 7:29 AM.
Operations terminated store 2859 at 7:35 AM.
At 8:31 AM, Fermilab experienced a network problem. The Feynman Computing Center lost its power. This problem was resolved later in the morning.
Operators prepared MiniBooNE for access at 9 AM.
While trying to set up a TeV 36x4 study-store, the TeV aborted due to a sector A4 quench. A quadrupole magnet (A4Q9) quenched. Experts replaced a quench card and installed a recorder to monitor the power supply.
At 1:13 PM, operators discovered at TeV quadrupole power supply (C:BQ0) with no output current. Earlier this morning operators noted this supply's odd behavior. Experts were called to investigate.
Pbar resumed stacking at 2 PM.
Operations put the TeV through a dry squeeze at 2:04 PM.
TeV power supply experts began replacing C:BQ0's reversing switch at 2:11 PM. They completed their work at 3:41 PM.
Operators prepared the Collider Detector Facility (CDF) for access 2:47 PM. The experimenters came out of access fifty minutes later.
Operations established beam to MiniBooNE at 3:29 PM.
Operations put the TeV through a dry squeeze at 4:37 PM and then at 6:44 PM established a 36x4 study-store.
At 6:45 PM, the TeV suffered a sector A4 quench.
A Booster RF station (BRF12) tripped off at 6:52 PM for the first of three times over an hour. It reset okay each time.
At 8:06 PM, the TeV sector A0 cryo compressor tripped off due to a loss of power, as did several others around the ring. The Duty Electrician reset the tripped breaker, which initially caused the problem. He could not find a cause for the trip.
Operator found the QPMs for TeV sector F3 & 4 not responding. They reloaded data. Operators believed that this was also due to the power outage.
Pbar resumed stacking at 10:16 PM.
Operations put the TeV through a dry squeeze at 11:47 PM.
Wednesday August 6
Operations put the TeV into shot setup at 1:24 AM. At 4:27 AM, Operations established store 2864 with an initial luminosity of 3.6E31. Pbar resumed stacking.
Operators tuned the Booster beam at 5:28 AM.
Comments and Suggestions
|last modified 8/6/2003 email Fermilab|